Living Space of Homes As reported by the Ghana Statistical Service. The mean living space for


single-family detached homes is 1742 sq. ft.


Assume a standard deviation of 568 sq. ft.


a. For samples of 25 single-family detached homes, determine the mean and standard deviation of the


variable


b. Repeat part (a) for a sample of size 500

1
Expert's answer
2022-03-25T03:18:17-0400

"a:\\\mu _{\bar{x}}=\mu =1742\\\sigma _{\bar{x}}=\frac{\sigma}{\sqrt{n}}=\frac{568}{\sqrt{25}}=113.6\\b:\\\mu _{\bar{x}}=\mu =1742\\\sigma _{\bar{x}}=\frac{\sigma}{\sqrt{n}}=\frac{568}{\sqrt{500}}=25.4017"